# CompAct: Compressed Activations for Memory-Efficient LLM Training

### v1

@@ 1 Introduction
Recent works tackle memory reductions by applying a low rank approximation to model parameters Hu et al. (2021); Lialin et al. (2023), or to gradients after the backward pass Muhamed et al. (2024); Hao et al. (2024); Zhao et al. (2024). However as seen in Figure 1 the main memory component is the computation graph itself. Its size also scales with batch size, in contrast with other memory components.
@@
In this work we introduce CompAct, a novel optimization technique that saves low-rank-compressed activations during the forward pass, instead of the full activation tensors. Consequently, the resulting gradients are low-rank as well, also reducing the size of optimizer states. As CompAct decompresses the gradients back to full size only for the update step, it compresses a large part of the compute graph, which in turn translates to major memory savings. Figure 2 presents an overview of our method, while Table 1 lists the scaling of different memory components, compared to other compression methods. CompAct is a logical next step from previous work, moving from low-rank parameters in Hu et al. (2021), through compressed low-rank gradients in Zhao et al. (2024), to compressed activations.
Overall, CompAct achieves savings of about 17.3% of peak device memory with practical batch sizes, for pretraining LLaMA-350M Touvron et al. (2023), and 50% for fine-tuning Roberta-Base Liu et al. (2019). As seen in Figure 1, CompAct’s memory reduction grows with model size. For LLaMA-65B, we can estimate a saving of 30% for LLaMA-65B. The same estimate for LLaMA-350M is 21%, which is 4% over the empirical value, giving a realistic estimate of 25%-30% memory saved for LLaMA-65B.
@@ 2 Related Work
Memory-efficient methods have become crucial for training LLMs, especially during pre-training, where the memory requirements for activations, gradients and optimizer states are significant as these models scale.
There are various approaches to making the training process more efficient, including mixed precision training Micikevicius et al. (2017) quantization methods Pan et al. (2022); Liu et al. (2022); Anonymous (2024b); Dettmers et al. (2023); Anonymous (2024a), parallelism techniques Dean et al. (2012); Li et al. (2014); Shoeybi et al. (2020); Huang et al. (2019) and low rank approximation methods. The latter being mostly unexplored for pretraining, it is the focus of this work.
@@
A recent approach, GaLore Zhao et al. (2024) utilizes the low-rank property of gradients to learn a full-rank model efficiently, even during pre-training. Instead of approximating the model’s weights, GaLore projects the gradients after the backward pass into a lower-dimensional subspace for the weight update, reducing the memory overhead without severely limiting the model’s expressiveness.
As GaLore relies on periodical SVD computations to maintain adequate low-rank approximations of the gradients, which is very expensive in both time and memory, a variety of works focus on relieving this computational cost, either by strategically choosing the projection matrices Anonymous (2024c), quantizing them Anonymous (2024b) or replacing them by random projections Muhamed et al. (2024); Hao et al. (2024). However, these methods remain inapplicable when pretraining large models, as peak device memory, is primarily determined by the activations stored on GPU (when the batch size is large) Anonymous (2024b); Muhamed et al. (2024).
Essentially, compared to GaLore, our approach may be viewed as a simple change in the order of operations, applying the compression one step before GaLore does, to the activations rather than to the gradients. As a result, CompAct satisfies their convergence theorem, which explains how it achieves comparable performance despite the drastic memory savings.
@@
Various works aim at reducing the memory cost of activations in deep learning in general. Yang et al. (2024) is a complimentary work focusing on saving activation memory generated by nonlinear functions and normalization layers, whereas our work focuses on the activations generated by linear layers. The two methods can be combined to achieve even greater savings, although some adaptation is required.
VeLoRA Miles et al. (2024) also aim to compress linear layer activations, however, they apply their paradigm to two specific layers of the model only, thus making their benefit marginal. We compare with their projection in our experimental section, see Section 4.4. In any case, both Miles et al. (2024) and Yang et al. (2024) remain unexplored for the setting of pretraining LLMs.
@@ 3.1 Background
(Convergence of GaLore with fixed projections). Suppose the gradient follows the parameteric form:
@@ 3.2 CompAct
An overview of the method is descibed in Algorithms 1,2,3.
@@ 4 Experiments
Model size 60M 130M 350M Perplexity GPU Peak Perplexity GPU Peak Perplexity GPU Peak Full-Rank 34.06 11.59 25.08 18.66 18.80 39.97 GaLore 34.88 11.56 25.36 18.48 18.95 39.24 CompAct $r=n/2$ 32.78 11.32 25.37 17.97 19.26 37.94 CompAct $r=n/4$ 34.41 10.80 26.98 16.78 20.45 34.71 CompAct $r=n/8$ 36.42 10.54 28.70 16.19 21.91 33.03 Training Tokens 1.1B 2.2B 6.4B
Table 2: Pretraining perplexity and peak GPU memory for different model sizes and different training techniques. Total training tokens are shown in the last row. As can be seen, CompAct reduces peak memory by up to 17% for LLaMA 350M, with comparable perplexity to baseline. For larger model sizes we estimate the total memory saving to be roughly 30%.
In this section, we evaluate CompAct on both pre-training (Section 4.1) and fine-tuning tasks (Section 4.2). In all experiments, we apply CompAct to all attention and MLP blocks in all layers of the model, except for the output projection in the attention mechanism. For further details, see Appendix A. Moreover, we provide a comparison of CompAct’s throughput and memory usage with other methods in Section 4.3, and explore various types of projection matrices in Section 4.4.
@@ 4.1 Pretraining
For pre-training, we apply CompAct to LLaMA-based models Touvron et al. (2023) of various sizes and train on the C4 (Colossal Clean Crawled Corpus) dataset, a commonly used dataset for training large-scale language models Raffel et al. (2023b). The models were trained without any data repetition.
Our experimental setup follows the methodology outlined in Zhao et al. (2024), using a LLaMA-based architecture that includes RMSNorm and SwiGLU activations Shazeer (2020). For each model size, we maintain the same set of hyperparameters, with the exception of the learning rate which was tuned.
In line with the approach proposed in GaLore, we apply a scaling factor of $\alpha=0.25$ to the CompAct layers. We also tuned the projection update, using $T=50$ period T after searching over [10,50,100,200,500$\infty$] for the LLaMA-60M model, which was then used consistently across all other model sizes. Further details regarding the training setup and hyperparameters can be found in Appendix A.1.
@@ 4.2 Finetuning
Table 4: Finetuning performance on several benchmarks for various compression rates with GaLore and CompAct. We report the empirical mean of three runs of our approach per task. Peak Memory was measured on RTE task. It is clear that both CompAct’s and GaLore’s performance is comparable with full finetuning, and very close to each other. However peak memory is vastly reduced with CompAct, with as much as 50% total memory saved. See Appendix B for the more details.
We fine-tune the pre-trained RoBERTa-base model Liu et al. (2019) on the GLUE benchmark, a widely used suite for evaluating NLP models across various tasks, including sentiment analysis and question answering Wang et al. (2019). We apply CompAct and compared its performance to GaLore. Following the training setup and hyperparameters from GaLore, we only tuned the learning rate. More details can be found in Appendix B
@@ 4.3 Peak Memory and Throughput
4.4 Ablation Projection Matrices
In this section, we explore the effects of different projection matrices on training performance when applied within the CompAct framework. We evaluate the following projection matrices:
@@ 4.4 Ablation Projection Matrices
As shown in Figure 4, using Gaussian projections with different seeds per layer slightly improved performance compared to a shared seed, suggesting that utilizing different subspaces for different layers enhances the learning capacity of the model. Additionally, the sparse JL projections performed comparably to the dense Gaussian projections. This is a promising result, suggesting the viability of efficient sparse operations to further improve the benefits of CompAct. Finally, incorporating the projection matrix from VeLoRa into CompAct performed poorly. This can be attributed to differences in how VeLoRa handles the backward pass, by projecting activations back and computing full-rank gradients. This gap is somewhat expected, as they only used their projection on two types of layers, whereas we applied the compression more broadly. For further details see Appedix C.
@@ 5 Conclusion
In this work, we presented CompAct, a memory-efficient method for training LLMs by compressing activations, gradients and optimizer states of linear layers. We demonstrate that CompAct achieves significant memory savings for training LLMs, reaching 25%-30% memory reduction for pretraining LLaMA-65B and 50% for Roberta-Base, with minimal impact on training throughput and performance. Our method is easily scalable, applicable to various model sizes, and is easily composable with other techniques.
By directly compressing the compute graph during training, CompAct targets a major component of peak device memory which was neglected in recent works. We believe this approach should guide future work for further memory gains. A good example could be incorporating sparse random projections into CompAct, which would reduce the computational cost associated with sampling and matrix operations. Another area for improvement is the approximation of intermediate activations, such as those generated by FlashAttention, Hadamard products and non-linearities, which require significant memory. By addressing these memory-intensive operations, CompAct’s memory reductions can be extended even further.
@@ 6 Limitations
Appendix A Appendix - Pretraining
This appendix supplies further details about our pretraining experiments.
A.1 Learning Rate
By searching for the optimal learning rate from the set $\{1e-3,5e-3,1e-2\}$, we found that the largest value $21e-$ 2 works best across all model sizes.
A.2 CompAct and FlashAttention
When measuring the memory consumption of our method we found that compressing the activation tensor of the output projection of the attention block does not result in memory savings, as the input tensor is still stored in memory due to the implementation of the FlashAttn mechanism Dao et al. (2022). Consequently, We chose to leave this layer uncompressed in our experiments, and leaving it for future work, where a custom cuda kernel for the flash attention should resolve this issue.
However, not compressing this layer introduced instability in the pretraining experiments due to the inconsistent learning rate across different linear layers induced by the scale $\alpha$. To mitigate this, we adjusted the learning rate specifically for this layer by a factor of 0.5.
A.3 Type Conversion in Normalization Layers
@@
Appendix B Appendix - Fine-tuning
To be comparable to the results reported in GaLore Zhao et al. (2024) as shown in Table 4 we report the same metrics as they did, namely, F1 score on QQP and MRPC, Matthew’s Correlation for CoLA, Pearson’s Correlation for STS-B, and classification accuracy for all others. The numbers reported for GaLore and Baseline are taken from Zhao et al. (2024). We report the average performance over three seeds due to the noisy behavior of the training process. All models were trained for 30 epochs with batch size 16, except for CoLA where we used batch size 32 as in GaLore, and a maximum sequence length of 512, a scale $\alpha=2$ was used with $r=8$ and $\alpha=4$ for $r=4$, all with $T=500$. Again, as in GaLore, we searched for a best learning rate per task, searching in [1e-5,2e-5,3e-5]. We apply the projection matrices or LoRA to target modules.
Appendix C Appendix - Projection Types

### v2

@@ 1 Introduction
Recent works tackle memory reductions by applying a low rank approximation to model parameters Hu et al. (2021); Lialin et al. (2023), or to gradients after the backward pass Muhamed et al. (2024); Hao et al. (2024); Zhao et al. (2024). However, as seen in Figure 1, the main memory component is the computation graph itself. Its size also scales with batch size, in contrast with other memory components.
@@
In this work, we introduce CompAct, a novel optimization technique that saves low-rank-compressed activations during the forward pass, instead of the full activation tensors. Consequently, the resulting gradients are low-rank as well, also reducing the size of optimizer states. As CompAct decompresses the gradients back to full size only for the update step, it compresses a large part of the compute graph, which in turn translates to major memory savings. Figure 2 presents an overview of our method, while Table 1 lists the scaling of different memory components, compared to other compression methods. CompAct is a logical next step from previous work, moving from low-rank parameters in Hu et al. (2021), through compressed low-rank gradients in Zhao et al. (2024), to compressed activations.
Overall, CompAct achieves savings of about 17.3% of peak device memory with practical batch sizes, for pretraining LLaMA-350M Touvron et al. (2023), and 50% for fine-tuning RoBERTa-Base Liu et al. (2019). As seen in Figure 1, the estimated memory reduction achieved by CompAct increases with model size. For LLaMA-65B, the estimated reduction is approximately 30%. For LLaMA-350M, the estimated reduction is around 21%, which is 4% higher than the observed empirical value. This suggests a realistic memory saving range of 25%-30% for LLaMA-65B.
@@ 2 Related Work
Memory-efficient methods have become crucial for training LLMs, especially during pretraining where the memory requirements for activations, gradients, and optimizer states are significant as these models scale.
There are various approaches to making the training process more efficient, including mixed precision training Micikevicius et al. (2017), quantization methods Pan et al. (2022); Liu et al. (2022); Anonymous (2024b); Dettmers et al. (2023); Anonymous (2024a), parallelism techniques Dean et al. (2012); Li et al. (2014); Shoeybi et al. (2020); Huang et al. (2019), and low rank approximation methods. The latter being mostly unexplored for pretraining, it is the focus of this work.
@@
A recent approach, GaLore Zhao et al. (2024) utilizes the low-rank property of gradients to learn a full-rank model efficiently, even during pretraining. Instead of approximating the model’s weights, GaLore projects the gradients after the backward pass into a lower-dimensional subspace for the weight update, reducing the memory overhead without severely limiting the model’s expressiveness.
As GaLore relies on periodic SVD computations to maintain adequate low-rank approximations of the gradients, which is very expensive in both time and memory, a variety of works focus on relieving this computational cost, either by strategically choosing the projection matrices Anonymous (2024c), quantizing them Anonymous (2024b), or replacing them by random projections Muhamed et al. (2024); Hao et al. (2024). However, these methods remain inapplicable when pretraining large models, as peak device memory is primarily determined by the activations stored on GPU (when the batch size is large) Anonymous (2024b); Muhamed et al. (2024).
Essentially, compared to GaLore, our approach may be viewed as a change in the order of operations, applying the compression one step before GaLore does: when storing activations in memory for the backward pass, rather than to the gradients when updating the optimizer state. As a result, CompAct satisfies their convergence theorem, which explains how it achieves comparable performance despite the drastic memory savings.
@@
Various works aim at reducing the memory cost of activations in deep learning in general. Yang et al. (2024) is a complementary work focusing on saving activation memory generated by nonlinear functions and normalization layers, whereas our work focuses on the activations generated by linear layers. The two methods can be combined to achieve even greater savings, although some adaptation is required.
VeLoRA Miles et al. (2024) also aims to compress linear layer activations, however, they apply their paradigm to two specific layers of the model only, thus making their benefit marginal. We compare with their projection in our experimental section, see Section 5. In any case, both Miles et al. (2024) and Yang et al. (2024) remain unexplored for the setting of pretraining LLMs.
@@ 3.1 Background
(Convergence of GaLore with fixed projections). Suppose the gradient follows the parametric form:
@@ 3.2 CompAct
An overview of the method is described in Algorithms 1,2,3.
@@ 4 Experiments
Model size 60M 130M 350M 1B Perplexity GPU Perplexity GPU Perplexity GPU Perplexity GPU Peak Peak Peak Peak Full-Rank 34.06 11.59 25.08 18.66 18.80 39.97 15.56 - GaLore 34.88 11.56 25.36 18.48 18.95 39.24 15.64 75.40 CompAct $r=n/2$ 32.78 11.32 25.37 17.97 19.26 37.94 17.40 72.82 CompAct $r=n/4$ 34.41 10.80 26.98 16.78 20.45 34.71 18.02 65.57 CompAct $r=n/8$ 36.42 10.54 28.70 16.19 21.91 33.03 19.23 61.88 Training Tokens 1.1B 2.2B 6.4B 13.1B
Table 2: Pretraining perplexity and peak GPU memory for different model sizes and different training techniques. Total training tokens are shown in the last row. As can be seen, CompAct reduces peak memory by up to 17% for LLaMA 350M, with comparable perplexity to baseline. For larger model sizes we estimate the total memory saving to be roughly 30%. The baseline for LLaMA 1B did not fit within the $\sim 81\,\text{GB}$ memory available at the same batch size.
In this section, we evaluate CompAct on both pretraining (Section 4.1) and finetuning tasks (Section 4.2). In all experiments, we apply CompAct to all attention and MLP blocks in all layers of the model, except for the output projection in the attention mechanism. For further details, see Appendix B. Moreover, we provide a comparison of CompAct’s throughput and memory usage with other methods in Section 4.3, and explore various types of projection matrices in Section 5.
@@ 4.1 Pretraining
For pretraining, we apply CompAct to LLaMA-based models Touvron et al. (2023) of various sizes and train on the C4 (Colossal Clean Crawled Corpus) dataset, a commonly used dataset for training large-scale language models Raffel et al. (2023b). The models were trained without any data repetition.
Our experimental setup follows the methodology outlined in Zhao et al. (2024), using a LLaMA-based architecture that includes RMSNorm and SwiGLU activations Shazeer (2020). For each model size, we maintain the same set of hyperparameters, with the exception of the learning rate and the projection update gap which were tuned. Further details regarding the training setup and hyperparameters can be found in Appendix B.
@@ 4.2 Finetuning
Table 4: Finetuning performance on several benchmarks for various compression rates with GaLore and CompAct. We report the empirical mean of three runs of our approach per task. Peak Memory was measured on RTE task. It is clear that both CompAct’s and GaLore’s performance is comparable with full finetuning, and very close to each other. However peak memory is vastly reduced with CompAct, with as much as 50% total memory saved. See Appendix C for the more details.
We finetune the pretrained RoBERTa-base model Liu et al. (2019) on the GLUE benchmark, a widely used suite for evaluating NLP models across various tasks, including sentiment analysis and question answering Wang et al. (2019). We apply CompAct and compared its performance to GaLore. Following the training setup and hyperparameters from GaLore, we only tuned the learning rate. More details can be found in Appendix C
@@ 4.3 Peak Memory and Throughput
5 Ablation Study
This section presents a series of ablation experiments examining how different design choices and hyperparameters affect training performance, memory usage, and convergence.
First, we explore the effects of different projection matrices on training performance when applied within the CompAct framework. We evaluate the following projection matrices:
@@ 5 Ablation Study
We opted not to experiment with Singular Value Decomposition (SVD)-based projections in our method due to practical considerations. More on SVD in Compact is discussed in Appendix A
@@
As shown in Figure 4, using Gaussian projections with different seeds per layer slightly improved performance compared to a shared seed, suggesting that utilizing different subspaces for different layers enhances the learning capacity of the model. Additionally, the sparse JL projections performed comparably to the dense Gaussian projections. This is a promising result, suggesting the viability of efficient sparse operations to further improve the benefits of CompAct. Finally, incorporating the projection matrix from VeLoRa into CompAct performed poorly. This can be attributed to differences in how VeLoRa handles the backward pass, by projecting activations back and computing full-rank gradients. This gap is somewhat expected, as they only used their projection on two types of layers, whereas we applied the compression more broadly. For the loss curves, see Appendix D.
Figure 5: Ablation on LlaMA 130M - Effect of varying projection update periods $T$ on performance across different ranks in CompAct.
Next, we examine how different training hyperparameters impact model convergence and memory efficiency.
Projection Update Period $T$:
We first analyze the influence of the projection update period $T$ on the convergence of of LLaMA-130M when trained with CompAct.To do so, we conduct experiments with varying update intervals. A learning rate of $35e-$ 3 is used by default, but if training becomes unstable at a given $T$, we reduce the learning rate until stability is achieved.
As shown in Figure 5, an optimal $T$ range emerges: updating the projection matrix too frequently or too infrequently both slow down convergence. This trend remains consistent across the couple of ranks we show.
Rank and Training Steps:
Next, we investigate the effect of rank and training duration on model convergence.Here, LLaMA-130M is trained with a projection update period of $T=200$ and an initial learning rate of $35e-$ 3, which is reduced to $34e-$ 3 if instability occurs.
Figure 6 shows the impact of varying the rank of the projection matrix over different training durations. As expected, lower ranks lead to more performance degradation, while increasing the rank improves model performance. Additionally, for any given rank, training for more steps yields better results.
Crucially, larger ranks require fewer training steps to match or surpass the baseline, whereas lower ranks need extended training to compensate for their reduced expressivity. This highlights a trade-off between memory, training time, and final performance. For instance, when constrained by a small GPU, using a more aggressive compression (e.g., rank 0.25) allows training to fit within memory limits. Depending on the number of training steps available, this trade-off can still yield competitive performance, and with sufficient steps, even exceed the baseline.
Figure 6: Perplexity vs. Rank - Effect of different ranks on the performance of CompAct-trained LLaMa-130M across varying training steps. The baseline (no compression) is trained for 20K steps.
Training Batch Size:
Finally, we examine how batch size affects peak GPU memory usage, demonstrating the benefits of CompAct when handling varying activation sizes. We measure the peak GPU memory consumption while training LLaMA-350M without gradient accumulation across different batch sizes. The comparison includes the baseline model, GaLore (with a rank of 256), and CompAct (with a rank of 0.25).
As shown in Figure 7, CompAct significantly reduces peak memory usage, and its benefit scales with batch size, whereas GaLore provides a fixed reduction in peak memory relative to the baseline. This difference arises because GaLore primarily compresses optimizer states, which are independent of batch size, while CompAct reduces the memory footprint of activations, which grow with batch size.
This result is particularly important because larger batch sizes are known to improve training throughputtempo. By lowering peak memory requirements, CompAct enables the use of larger batches, potentially increasing training efficiency without exceeding hardware constraints.
Figure 7: Peak GPU memory vs. Batch Size. Peak GPU memory usage of LLaMa-350M for different batch sizes is shown for CompAct, GaLore, and the baseline.
@@ 6 Conclusion
In this work, we presented CompAct, a memory-efficient method for training LLMs by compressing activations, gradients, and optimizer states of linear layers. We demonstrate that CompAct achieves significant memory savings for training LLMs, reaching 25%-30% memory reduction for pretraining LLaMA-65B and 50% for RoBERTa-Base, with minimal impact on training throughput and performance. Our method is easily scalable, applicable to various model sizes, and is easily composable with other techniques.
By directly compressing the compute graph during training, CompAct targets a major component of peak device memory which was neglected in recent works. We believe this approach should guide future work for further memory gains. A good example could be incorporating sparse random projections into CompAct, which would reduce the computational cost associated with sampling and matrix operations. Another area for improvement is the approximation of intermediate activations, such as those generated by FlashAttention, Hadamard products, and non-linearities, which require significant memory. By addressing these memory-intensive operations, CompAct’s memory reductions can be extended even further.
@@ 7 Limitations
Appendix A CompAct with SVD
In this work, random projections are used to compress activations, thereby avoiding the costly SVD step. Performing an SVD on the activation tensor $x\in\mathbb{R}^{b\times l\times n}$, where $b$ is the batch size and $l$ is the sequence length, can introduce considerable overhead since $b\times l$ is typically large. By contrast, GaLore applies SVD to the gradient $G\in\mathbb{R}^{n\times m}$, which does not depend on $b$ or $l$, making it generally less expensive than compressing activations via SVD.
Nonetheless, one could adopt GaLore’s SVD-based projection within the CompAct framework by running an iteration with uncompressed activations to compute the gradient, from which the SVD projection is derived and then stored for subsequent updates. However, each time the projection matrix is updated, storing the entire activation tensor would significantly increase the GPU memory peak, since the activations typically dominate memory usage. A potential mitigation strategy might involve updating the projection matrix on smaller batches to reduce peak memory requirements. Further exploration of this approach is left for future work.
Appendix B Pretraining
This appendix provides further details about our pre-training experiments.
B.1 Hyperparameters
We adopt the training setup outlined in Zhao et al. (2024), and apply compact to LLaMA-based models of various sizes. Table 5 outlines the amount of data and steps used to train the models.
Model Size Steps Training Tokens 60M 10K 1.3 B 130M 20K 2.6 B 350M 60K 7.8 B 1B 100K 13.1 B
Table 5: Training Step for Llama models.
For all the trained models, we use a maximum sequence length of 256, with a batch size of 512. We tune the optimal learning rate for each model size from the range $21e-3\leq lr\leq 1e-$ 2, and choose the best one based on the validation perplexity. We also adopt a learning rate warmup for the first 10% of the training steps and use a cosine learning rate scheduler that decreases to 0.1 of the initial learning rate.
We use a constant scaling factor of $\alpha=0.25$. Additionally, we tuned the projection update, using $T=50$ period T for after searching over [1,10,50,100,200,500$\infty$] for the LLaMA-130M model, which was then used to train 60M - 350M. For the 1B model, we use $T=200$.
B.2 CompAct and FlashAttention
We applied CompAct to all linear modules within each Transformer block of our LLaMA-based models, except for the output-projection layer in the attention mechanism. FlashAttention Dao et al. (2022), a fast and efficient implementation of attention, is widely adopted in many architectures, including the models used in this work for both pretraining and finetuning.
However, FlashAttention stores its output in memory as part of the activation memory. This same tensor is then fed into the output-projection layer of the attention block, meaning that these two layers share the same activation memory. Consequently, compressing the activation tensor of the attention’s output-projection does not result in overall memory savings, because the FlashAttention output tensor remains in memory. Further, compressing the shared activation between these two layers would introduce errors in the gradient with respect to the FlashAttention input, potentially causing error accumulation across layers. Addressing this issue lies outside the current scope of CompAct.
As a result, we left the output-projection layer uncompressed in our experiments. In future work, a custom CUDA kernel for FlashAttention may allow compression of this layer without introducing errors.
However, not compressing this layer introduced instability in the pretraining experiments when training with larger learning rates, due to the inconsistent learning rate across different linear layers induced by the scale $\alpha$. To mitigate this, one can either adjust the learning rate of the output-projection layer by a scale $\alpha_{out}$, or simply use a smaller learning rate.
In our experiments, when training the smaller models 60M-350M we used a large learning rate of $0.01$ and scaled the learning of the output projection with $\alpha_{out}=0.5\alpha$, while for the larger model we simply used a smaller learning rate of $0.003$.
B.3 Type Conversion in Normalization Layers
@@
Appendix C Fine-tuning
To be comparable to the results reported in GaLore Zhao et al. (2024) as shown in Table 4 we report the same metrics as they did, namely, F1 score on QQP and MRPC, Matthew’s Correlation for CoLA, Pearson’s Correlation for STS-B, and classification accuracy for all others. The numbers reported for GaLore and Baseline are taken from Zhao et al. (2024). We report the average performance over three seeds due to the noisy behavior of the training process. All models were trained for 30 epochs with batch size 16, except for CoLA where we used batch size 32 as in GaLore, and a maximum sequence length of 512, a scale $\alpha=2$ was used with $r=8$ and $\alpha=4$ for $r=4$, all with $T=500$. Again, as in GaLore, we searched for a best learning rate per task, searching in $\{1e-5,2e-5,3e-5\}$.
Appendix D Choice of Projection Matrix - Loss Curve